OpenStudy (countryalltheway):

Can some one please help me with these please i am willing to fan and medal

OpenStudy (countryalltheway):

x - 2y = 3 5x + 3y = 2 The lines whose equations are shown intersect at (1, -1) (-1, 1) (0, 3/2) 5x + 2y = 6 3x + y = 4 Which of the following is part of the solution to the system of equations? x = -2 y = -2 x = 1 Solve the system by the elimination method. x + y - 6 = 0 x - y - 8 = 0 When you eliminate y , what is the resulting equation? 2x = -14 2x = 14 -2x = 14

OpenStudy (countryalltheway):

@StudyGurl14 please help

OpenStudy (countryalltheway):

please i am going insane because i am stuck on these

OpenStudy (studygurl14):

Only one question at a time please. I'll help you with the first question: x - 2y = 3 5x + 3y = 2 The lines whose equations are shown intersect at (1, -1) (-1, 1) (0, 3/2) The easiest and fastest way to do this is to just graph these two equations at https://www.desmos.com/calculator

OpenStudy (countryalltheway):

@StudyGurl14 Okay so how do i do it

OpenStudy (countryalltheway):

@Data_LG2 Hey can u please please help me

OpenStudy (studygurl14):

Just type your equations in on the place to the left.

OpenStudy (countryalltheway):

Please @Data_LG2

OpenStudy (countryalltheway):

@Nnesha

OpenStudy (countryalltheway):

i got the first one done i dont get the rest of them

OpenStudy (studygurl14):

You got #1? @Countryalltheway

OpenStudy (countryalltheway):

yah

OpenStudy (countryalltheway):

its A right

OpenStudy (countryalltheway):

i dont get the next two

OpenStudy (studygurl14):

Correct. nice job. :)

OpenStudy (studygurl14):

You're not supposed to put more than one question per thread. Since you have more questions, close this thread and then open the next one in a new thread.

OpenStudy (countryalltheway):

these are the rest i need help with

OpenStudy (countryalltheway):

1.) 2x + y = 3 2.) x - 2y = -1 If equation 1 is multiplied by 2 and then the equations are added, the result is 3x = 2 3x = 5 5x = 5 x + y = k 2x + 3y = k + 1 The point of intersection of the lines has an x-coordinate of 2k + 1 -2k + 1 2k - 1 Solve the system by the elimination method. Check your work. x - 3y = 0 3y - 6 = 2x {(-6, -1/2)} {(-6, -2)} {(-2, -6)}

OpenStudy (countryalltheway):

really that will take a lot of threads

OpenStudy (studygurl14):

Yes really. Reason number one: multiple questions creates long threads that causes openstudy to lag #2: what if one person helps you on one question, and then has to leave and someone else helps you on another? You can only give one medal per thread.